1. The price of a pedicure is $40.

At an old income of $120,000, Jaime’s yearly demand for pedicures was Qd = 50 – 0.25P

At a new income of $150,000, Jaime’s yearly demand for pedicures is Qd = 80 – 0.25P

Jaime’s income elasticity of demand for pedicures is __________.

(Your answer may be a decimal. If so, round to the nearest tenth)
Answer #1

Ans: 2.4

Explanation:

At the income of $120,000,

Qd = 50 - 0.25P = 50 - 0.25(40) = 40

At the income of $150,000,

Qd = 80 - 0.25P = 80 - 0.25(40) = 70

So, here Y1 = 120,000    Q1 = 40

               Y2 = 150,000    Q2 = 70

Income elasticity of demand for pedicures = (Q2 - Q1) / (Y2 - Y1) * (Y1 + Y2) / (Q1 + Q2)

                                                                   = (70 - 40) / (150,000 - 120,000) * (120,000 + 150,000) / (40 + 70)

                                                                   = (30 / 30,000) * (270,000 / 110)

                                                                   = 8,100,000 / 3,300,000

                                                                   = 2.4